OpenStudy (afrodiddle):

What is the exact value of the expression the square root 72. - the square root of 8. + the square root of 128.? Simplify if possible.

OpenStudy (afrodiddle):

I think I know it

OpenStudy (afrodiddle):

am I right?

OpenStudy (afrodiddle):

\[12\sqrt{2}\]

OpenStudy (welshfella):

factor the numbers first - try and get a perfect square in the factors eg sqrt 72 = sqrt 36 * sqrt 2 = 6 sqrt 2

OpenStudy (welshfella):

sqrt 8 = sqrt 4 * sqrt 2 can you simplify sqrt 128?

OpenStudy (here_to_help15):

\[\sqrt{72}-\sqrt{8}+\sqrt{128}\] this , right ? can you factor, 72 =.... ? 8=...? 128 =...?

OpenStudy (afrodiddle):

72 is 6 square roof of 2 8 is 2 square root of 2 128 is 8 square root of 2

OpenStudy (here_to_help15):

good! now just factor out sqrt 2 from each term :) can u ?

OpenStudy (afrodiddle):

whaa?

OpenStudy (welshfella):

6 sqrt2 - 2 sqrt2 + 8 sqrt2

OpenStudy (welshfella):

= ?

OpenStudy (afrodiddle):

Ahhh. So it is 12 square root of 2?

OpenStudy (welshfella):

yes

OpenStudy (afrodiddle):

I did that earlier, I just got confused on what he/she said. >.<

OpenStudy (welshfella):

yes but you wrote it 3 times???

OpenStudy (here_to_help15):

Sorry

OpenStudy (afrodiddle):

What do you mean?

OpenStudy (welshfella):

you wrote 12 sqrt2 3 times

OpenStudy (afrodiddle):

Where?

OpenStudy (welshfella):

in your 3rd post

OpenStudy (afrodiddle):

It only shows up once for me.

OpenStudy (welshfella):

well thats weird - the first 12 sqrt2 shows up in large font and the next 2 in a smaller font - that so strange.....

OpenStudy (afrodiddle):

I am haunted :O

OpenStudy (welshfella):

lol